calculate the volume that will be occupied by 20.0g carbon dioxide at 25 degree celcius and 1.25 atm.


1
Expert's answer
2022-03-12T01:59:19-0500

Mass of carbon dioxide = 20.0 g ,

Mole of carbon dioxide = "\frac{20}{44}" = 0.45

pressure = 1.25 atm , temp = 25 oC = 298 K

PV=nRT

1.25×V= 0.45×0.0821×298

V= 0.11 L